# JEPA4Rec: Learning Effective Language Representations for Sequential Recommendation via Joint Embedding Predictive Architecture

### v1

@@ 4.3 Efficient Learning Representation Performance
Figure 5: NDCG@10 comparision between models over different size of training data
Low-Resource Training We compare ID-based models (SASRec, LlamaRec) with text-only models (RecFormer, JEPA4Rec) across varying data ratios. Figure 5 shows text-only models consistently outperform ID-based ones, especially with just $1\%$ or $5\%$ of the data. With limited data, ID-based models struggle as unseen items get random embeddings, while text-based models leverage prior knowledge. As data increases, SASRec and LlamaRec improve rapidly. Despite this improvement, these results confirm that language-based models excel in low-resource settings by utilizing item text for better recommendations.
@@ 4.4 Ablation Study
We perform an ablation study on Scientific (cross-domain) and Online Retail (cross-platform) datasets to assess JEPA4Rec’s key components, detailed results in Table 3. (1) w/o MLM loss: Removing Masked Language Modeling (MLM) reduces performance across all metrics, highlighting its role in bridging the semantic gap between Longformer’s pre-trained knowledge and the recommendation dataset; (2) w/o pre-training: Performance drops significantly, emphasizing the necessity of pre-training for generalizable item embeddings and common-sense user preference learning; (3) w/o token type embeddings: While it has little impact on Scientific, it significantly lowers Recall@10 on Online Retail, showing its importance in distinguishing patterns within item sentences.
@@ Appendix A Joint Embedding Predictive Architecture
Appendix B Baselines
@@
  • Text-only Methods: UniSRec (Hou et al., 2022) employs text-based item representations from a pre-trained language model, adapting to new domains through an MoE-enhanced adaptor. VQ-Rec (Hou et al., 2023) mitigates the over-reliance on textual features in transferable recommenders by mapping item text to discrete codes, which are then used to retrieve item representations from a code embedding table (text $\rightarrow$ code $\rightarrow$ representation). We initialize them with pre-trained parameters provided by the authors and finetune them on target domains. RecFormer (Li et al., 2023a) introduces a framework that formulates items as text sequences and employs a bidirectional Transformer to learn language representations for sequential recommendation.
Appendix C Experiment Settings
@@
Appendix D Case Study
(a) The purchase history of a user in Musical Instruments dataset. Comparison of recommended item before and after revealing information. The revealed information consists of two words: ”Musical” and ”Amplifier”.
@@ Appendix D Case Study
We analyze how text-only models adapt with partial next-item information by revealing different word ratios from item sentences in a zero-shot setting on the Instruments dataset, detailed in Figure 7. Both JEPA4Rec and RecFormer improve Recall@10, but JEPA4Rec shows a larger gain due to its mapping loss pre-training. Notably, JEPA4Rec’s performance doubles with just $10\%$ of item information, demonstrating its ability to leverage minimal textual cues. For a user interested in music production and guitar-related items, JEPA4Rec ranked the ground-truth item 10th initially but moved it to 1st after revealing partial information. This highlights its strong adaptability, making it well-suited for real-world scenarios where only limited item context is available.
Appendix E More zero-shot Experiments
@@
We conducted a comparison of JEPA4Rec’s zero-shot capability when pre-trained on three Amazon datasets against the official public checkpoint of RecFormer which was pre-trained on seven Amazon datasets. We can observe that the recommendation capabilities of the two models are quite similar, demonstrating JEPA4Rec’s robust training ability that does not depend on large-scale data.
Appendix F More ablation studies
(a) Scientific
(b) Office
Figure 8: Effect of Masking Ratio on Performance
We wanted to determine the appropriate masking ratio for next item prediction, so we experimented with masking different ratios of next item sentences. We pre-trained JEPA4Rec from scratch on the Scientific (small-scale) and Office (large-scale) dataset. The average results after 5 times in Figure 8 show that the optimal ratio is masking $50\%$ of the tokens in next item sentences.

### v3

@@ 4.3 Efficient Learning Representation Performance
Figure 5: NDCG@10 comparison between models over different sizes of training data
Low-Resource Training We compare ID-based models (SASRec, LlamaRec) with text-only models (RecFormer, JEPA4Rec) across varying training data ratios on both the Instruments and Scientific datasets. As shown in Figure 5, text-only models consistently outperform ID-based ones, particularly under low-resource conditions (e.g., $1\%$ or $5\%$ of the data). This performance gap is attributed to the ability of language-based models to utilize semantic item descriptions during pre-training, enabling better generalization to unseen items. In contrast, ID-based models assign random embeddings to new items, making it difficult to generate meaningful recommendations with sparse data. As the training data increases, ID-based models such as SASRec and LlamaRec show noticeable improvements, reducing the performance gap. Nevertheless, JEPA4Rec remains the top-performing model across all data scales, highlighting the robustness of contrastive language modeling for recommendation.
@@ 4.4.1 Study of JEPA4Rec
We perform an ablation study on Scientific (cross-domain) and Online Retail (cross-platform) datasets to assess JEPA4Rec’s key components, with detailed results in Table 3. (1) w/o MLM loss: Removing Masked Language Modeling (MLM) reduces performance across all metrics, highlighting its role in bridging the semantic gap between Longformer’s pre-trained knowledge and the recommendation dataset; (2) w/o pre-training: Performance drops significantly, emphasizing the necessity of pre-training for generalizable item embeddings and common-sense user preference learning; (3) w/o token type embeddings: While it has little impact on Scientific, it significantly lowers Recall@10 on Online Retail, showing its importance in distinguishing patterns within item sentences.
4.4.2 Masking Ratio Analysis
(a) Scientific
(b) Office
Figure 6: Effect of Masking Ratio on Performance
To investigate the effect of varying masking ratios on next-item prediction performance, we conducted a comprehensive ablation study by masking different proportions of tokens in next-item sentences during pre-training. Specifically, we trained our JEPA4Rec model from scratch on both the Scientific dataset (representing small-scale data) and the Office dataset (representing large-scale data). The results, averaged over five runs, are presented in Figure 6. We observe a consistent pattern across both datasets: the NDCG@10 score improves as the masking ratio increases from $10\%$ to $50\%$, and then degrades when further increased to $80\%$. This trend suggests that moderate masking (around $50\%$) provides the most effective learning signal for the model, balancing information retention with learning generalizable representations.
4.4.3 Additional Experimental Results.
Further experiments are included in Appendix G, covering case studies, robustness analysis, and additional ablation. First, we examine JEPA4Rec’s zero-shot capability and adaptability to partial item information. Case studies show that JEPA4Rec can accurately rank relevant items with minimal text cues, significantly outperforming RecFormer. We also provide extended zero-shot evaluations across six datasets, where JEPA4Rec remains competitive despite being pre-trained on fewer domains. Lastly, ablation studies on token embedding components and loss functions demonstrate the importance of MLM loss, token structure embeddings, and contrastive learning, each contributing significantly to model effectiveness.
@@ Appendix A Joint Embedding Predictive Architecture
Appendix B Details of the Tokenizer Component
Figure 8: Tokenizer Component
Figure 8 illustrates the Tokenizer Component in detail. This component is responsible for transforming item metadata into sequence representations through a masking strategy that simulates partial observation. Specifically, it flattens item metadata and applies token-level masking to create a masked input sequence, which is then passed through the Context Encoder. Simultaneously, the complete (unmasked) input is processed by the Target Encoder. These parallel encoding paths enable the model to learn robust representations by contrasting partial and full information.
In detail, to encode item sentences, we follow previous work in Sun et al. (2019); Geng et al. (2022) to construct four different types of embeddings:
Appendix C Baselines
@@
  • Text-only Methods: UniSRec (Hou et al., 2022) employs text-based item representations from a pre-trained language model, adapting to new domains through an MoE-enhanced adaptor. VQ-Rec (Hou et al., 2023) mitigates the over-reliance on textual features in transferable recommenders by mapping item text to discrete codes, which are then used to retrieve item representations from a code embedding table (text $\rightarrow$ code $\rightarrow$ representation). We initialize them with pre-trained parameters provided by the authors and fine-tune them on target domains. RecFormer (Li et al., 2023a) introduces a framework that formulates items as text sequences and employs a bidirectional Transformer to learn language representations for sequential recommendation.
Appendix D Experiment Settings
@@
Appendix E Case Study
(a) The purchase history of a user in Musical Instruments dataset. Comparison of the recommended item before and after revealing information. The revealed information consists of two words: ”Musical” and ”Amplifier”.
@@ Appendix E Case Study
Firt case study, we analyze how text-only models adapt with partial next-item information by revealing different word ratios from item sentences in a zero-shot setting on the Instruments dataset, detailed in Figure 9. Both JEPA4Rec and RecFormer improve Recall@10, but JEPA4Rec shows a larger gain due to its mapping loss pre-training. Notably, JEPA4Rec’s performance doubles with just $10\%$ of item information, demonstrating its ability to leverage minimal textual cues. For a user interested in music production and guitar-related items, JEPA4Rec ranked the ground-truth item 10th initially but moved it to 1st after revealing partial information. This highlights its strong adaptability, making it well-suited for real-world scenarios where only limited item context is available.
Second case study, to evaluate the robustness of our approach when item information is incomplete, we conducted a systematic evaluation by randomly dropping 20%, 40%, and 60% of tokens from item sequences at the target domain during the fine-tuning process. This simulates real-world scenarios where item descriptions may be incomplete or partially available. The masking is applied to the flattened item sentences containing “Title … Brand … Category …” information.
Masking Ratio Method Scientific (NDCG@10) Instrument (NDCG@10) Office (NDCG@10)
Full info RecFormer 0.1198 0.1001 0.1151
JEPA4Rec 0.1282 0.1057 0.1276
20% masked RecFormer 0.1145 0.0971 0.1207
JEPA4Rec 0.1223 0.0999 0.1242
40% masked RecFormer 0.1114 0.0954 0.1174
JEPA4Rec 0.1204 0.0997 0.1233
60% masked RecFormer 0.1088 0.0905 0.1093
JEPA4Rec 0.1201 0.0979 0.1197
Table 4: Robustness evaluation with different masking ratios
These results demonstrate JEPA4Rec’s superior robustness when making recommendations based on partial item information. Notably, JEPA4Rec maintains stable performance even when 60% of item tokens are masked, with only modest performance degradation (e.g., from 0.1282 to 0.1201 on the Scientific dataset). In contrast, RecFormer shows a significant accuracy downgrade as the masking ratio increases, with performance declining more substantially (e.g., from 0.1198 to 0.1088 on the Scientific dataset). This robustness stems from JEPA4Rec’s ability to learn item representations in the embedding space through its mapping loss pre-training, enabling effective inference even with incomplete textual descriptions. JEPA4Rec consistently outperforms RecFormer by $\sim 3-5\%$ on the NDCG metric across all masking scenarios.
Appendix F More zero-shot Experiments
@@
We conducted a comparison of JEPA4Rec’s zero-shot capability when pre-trained on three Amazon datasets against the official public checkpoint of RecFormer, which was pre-trained on seven Amazon datasets. We can observe that the recommendation capabilities of the two models are quite similar, demonstrating JEPA4Rec’s robust training ability that does not depend on large-scale data.
Appendix G More ablation studies
G.1 Token Embedding Component Analysis
We evaluate the importance of different token embedding types:
Experiment 1: Removing MLM Loss We remove the MLM loss component to assess how much semantic information learning from textual data contributes to JEPA4Rec’s recommendation performance.
Experiment 2: Removing Token Type Embedding ($\mathcal{C}$) This experiment evaluates the importance of token type embedding in distinguishing different types of textual information within item representations.
Experiment 3: Removing Token Position Embedding ($\mathcal{B}$) and MLM Loss Token position embedding $\mathcal{B}$ is crucial for language models to determine token positions. We remove both $\mathcal{B}$ and MLM loss to assess JEPA4Rec’s recommendation capability when losing components that help learn semantic information from e-commerce text data.
Experiment 4: Removing MLM Loss, Token Position Embedding ($\mathcal{B}$), and Token Type Embedding ($\mathcal{C}$) This experiment determines JEPA4Rec’s performance when relying solely on item representation recovery through L2 Mapping loss.
Note on Item Position Embedding ($\mathcal{D}$): We do not remove $\mathcal{D}$ as it is essential for determining item positions in history sequences and enabling JEPA4Rec to recover complete item representations.
Variant Scientific Online Retail
NDCG@10 Recall@10 MRR NDCG@10 Recall@10 MRR
Full JEPA4Rec 0.1282 0.1761 0.1190 0.1266 0.2429 0.0985
w/o MLM loss 0.1170 0.1653 0.1128 0.1118 0.2216 0.0858
w/o Token Type ($\mathcal{C}$) 0.1251 0.1749 0.1072 0.1231 0.1755 0.0915
w/o MLM + Token Pos ($\mathcal{B}$) 0.1102 0.1542 0.1021 0.1094 0.1898 0.0824
w/o MLM + $\mathcal{B}$ + Token Type ($\mathcal{C}$) 0.0967 0.1398 0.0889 0.1019 0.1687 0.0792
Table 6: Token embedding component analysis
These results demonstrate:
  1. 1.
    MLM loss is crucial for performance as it helps the model learn additional semantic information from the textual data of items. The performance drop when removing MLM loss (from 0.1282 to 0.1170 NDCG@10 on the Scientific dataset) shows its significant contribution to understanding item semantics.
  2. 2.
    All embedding components are important for the model’s effectiveness:
    • Token type embedding ($\mathcal{C}$) provides discriminative power for different attribute patterns, with noticeable performance degradation when removed.
    • Token position embedding ($\mathcal{B}$) is essential for understanding token sequence structure.
    • The combined removal of multiple components leads to progressively worse performance, confirming that each embedding type contributes unique and valuable information to the model.
G.2 Ablation Studies on Loss Components
We conduct ablation studies on different loss components to understand their contributions:
Experiment 1: Pre-training when dropping contrastive loss We evaluate JEPA4Rec’s ability to learn item representations in the embedding space by using only the Mapping Loss and MLM during pre-training. This experiment demonstrates how well the item representation recovery mechanism alone contributes to recommendation performance without the contrastive learning component.
Experiment 2: Fine-tuning with BPR Loss During fine-tuning, we replace the contrastive loss with the well-known Bayesian Personalized Ranking (BPR) loss to assess the impact of different ranking objectives:
$\mathcal{L}_{BPR}=-\sum_{(u,i,j)\in D_{S}}\ln\sigma(\hat{r}_{ui}-\hat{r}_{uj})$
where $D_{S}=\{(u,i,j)|i\in I_{u}^{+}\land j\in I\setminus I_{u}^{+}\}$ represents the training data with positive item $i$ and one sampled negative item $j$ for user $u$, and $\hat{r}_{ui}$ denotes the predicted preference score.
Variant Scientific Online Retail
NDCG@10 Recall@10 MRR NDCG@10 Recall@10 MRR
Full JEPA4Rec 0.1282 0.1761 0.1190 0.1266 0.2429 0.0985
Pre-training: Drop contrastive loss 0.1216 0.1634 0.1087 0.1064 0.1853 0.0892
Fine-tuning: BPR Loss 0.1245 0.1723 0.1156 0.0934 0.1637 0.0865
Table 7: Ablation studies on loss components
These results demonstrate that contrastive loss with its ability to sample multiple negative items remains well-suited for JEPA4Rec. The performance degradation when replacing contrastive loss with BPR loss (which samples only one negative item per positive) shows that the multi-negative sampling strategy of contrastive learning provides richer training signals. This is particularly evident in the Online Retail dataset, where BPR loss shows significant performance drops across all metrics, highlighting the importance of diverse negative sampling for effective representation learning in our framework.